OpenStudy (anonymous):

the perimeter of a rectangle is 116 if side WX=2x+3 and side VW equal 3x find the length of WX without variables

OpenStudy (anonymous):

perimeter = 2* ( A + b ) =116 2* (2X +3 + 3X ) = 116 5X+3 = 58 X= 11

OpenStudy (anonymous):

sides are 2x + 3 = 25 and 3x = 33

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

VW